Accelerator Update for May 13, 2009 - May 15, 2009 |
Wednesday May 13 The day shift began with Operations monitoring the Tevatron (TeV) store 7044 (109.86E30) and the Recycler stash (183.7E10), with the Antiproton Source stacking (18.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations disabled the MTest beam at 12:35 PM, at which point a power supply LCW leak on MT2WD2 took down supplies in M01 through M03. Investigation and repair lasted through to the end of the day shift. Operations terminated store 7044 at 5:11 PM and then put the TeV in shot setup. TeV personnel conducted coupling studies during shot setup. Operations established store 7045 at 7:46 PM with an average initial luminosity of 324.86E30. The MTest LCW leak was repaired early on the evening shift, but not before MTest was scheduled to shut off. Operations transferred antiprotons to the Recycler as appropriate. Thursday May 14 The midnight shift began with Operations monitoring store 7045 (160.1E30) and the stash (142.7E10), with Pbar stacking (23.3E10), and with NuMI and MiniBooNE taking beam. Operations reported that a Linac RF station (LRF4) had caused downtime due to problems with its power amplifier. Also, a MI RF station (MIRF13) kept tripping off. Operations established beam to MTest at 2 AM. Operations disabled the NuMI beam at 7 AM in preparation for a scheduled access to repair a dehumidifier. The MTest beam unexpectedly went down at 8:40 AM due to scheduled Meson area power outages. Beam returned to the experiment once a power workaround was put in place. NuMI personnel accessed MI-65 at 9:45 AM. Operations terminated store 7045 at 10:16 AM. D0 accessed their collision hall. Operations had the TeV in shot setup by 2:30 PM. There were problems with Recycler losing antiprotons when a corrector chassis tripped off. Once things were stabilized, shot setup resumed, but ended abruptly in a quench just prior to ramping. Operators and TeV personnel conducted a quench investigation. Accesses were made to F-Sector, CDF, and D0. Several minor repair jobs in the TeV were completed as well. Linac personnel began work on LRF4 during the TeV accesses. NuMI came out of MI-65 around 2:42 PM, but a problem held their beam off until near the end of the day shift. Operations put the TeV through a dry squeeze at 10:15 PM and remained there through the rest of the evening shift. Operations transferred antiprotons to the Recycler as appropriate. Friday May 15 The midnight shift began with the TeV in a dry squeeze, with Operations monitoring the stash (82.6E10), with Pbar stacking (19.9E10), and with NuMI and MiniBooNE beam off due to the LRF4 work. Problems with a Recycler corrector bulk supply required EE Support personnel intervention. No beam was lost. LRF4 came back up shortly after midnight. The TeV dry squeeze was followed by several wet squeezes and orbit corrections. Operations established beam to MTest at 2 AM. Operations put the TeV into shot setup at 4:14 AM. Operations established store 7054 at 6:07 AM with an average initial luminosity of 181.18E30. Operations transferred antiprotons to the Recycler as appropriate. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/15/2009 email Fermilab |
| Security, Privacy, Legal | |